OpenStudy (anonymous):

please help!:) which of the following is a solution for the system y=-3x+5 and y=2x+5?

OpenStudy (radar):

y=y 3x+5=2x+5, x=5-5 x=0 y=5

OpenStudy (anonymous):

so 0,5?

OpenStudy (radar):

Yes